# Pixel-Perfect Depth with Semantics-Prompted Diffusion Transformers

### v1

@@ 2.1 Monocular Depth Estimation
Depth estimation can be broadly categorized into monocular yang2024depthv2 ; wang2025moge , stereo igev ; xu2023accurate ; igevpp ; cheng2025monster , and sparse depth completion promptda methods. Early monocular depth estimation methods relied primarily on manually designed features saxena2008make3d ; hoiem2007recovering . The advent of neural networks revolutionized the field, though initial approaches eigen2014depth ; eigen2015predicting struggled with cross-dataset generalization. To address this limitation, scale-invariant and relative loss midas are introduced, enabling multi-dataset Li2018:CVPR ; blendedmvs ; diml ; hrwsi ; tartanair ; irs ; wsvd ; redweb ; hypersim ; liang2025sood++ training. Recent methods focus on improving the generalization ability yang2024depthv2 ; depthpro , depth consistency yang2024depthanyvideo ; chen2025video ; hu2024depthcrafter ; ke2024video , and metric scale bhat2023zoedepth ; li2024patchfusion ; li2024patchrefiner ; yin2023metric3d ; guizilini2023towards ; leres ; hu2024metric3dv2 ; piccinelli2024unidepth ; promptda of depth estimation. These methods converge towards using transformer-based architectures dpt . Concurrent works wang2025moge ; moge2 ; xu2025geometrycrafter explore point cloud representations to improve depth estimation performance. Several recent methods ji2023ddp ; duan2024diffusiondepth ; saxena2023monocular ; saxena2023surprising ; saxena2023zero ; zhao2023unleashing have attempted to use diffusion models for metric depth estimation. In contrast, our method focuses on relative depth and demonstrates improved generalization and fine-grained detail across a wide range of real-world scenes. Furthermore, our model significantly differs from these methods by introducing Semantics-Prompted DiT, which incorporates pretrained high-level semantic representations into the diffusion process, greatly enhancing performance.
@@ 2.2 Diffusion Generative Models
Diffusion generative models ddpm ; ddim ; dit ; yu2024representation ; yao2024fasterdit ; yao2025reconstruction have demonstrated impressive results in image and video generation. Early approaches ddpm ; ho2022classifier ; ho2022cascaded such as DDPM ddpm operate directly in the pixel space, enabling high-fidelity generation but incurring significant computational costs, especially at high resolutions. To address this limitation, Latent Diffusion Models perform the diffusion process in a lower-dimensional latent space obtained via a VAE, as popularized by Stable Diffusion stablediff . This design significantly improves training and inference efficiency and has been widely adopted in recent works sd3 ; yao2025reconstruction ; yu2024representation ; zhu2024dig ; flux2024 ; podell2023sdxl ; yang2024cogvideox .
@@ 4.2 Ablations and Analysis
Component-wise ablation analysis. We adopt the DiT dit model as our baseline and conduct ablations on our proposed modules. Quantitative results are shown in Table 2. Directly performing diffusion generation in high-resolution pixel space is highly challenging due to substantial computational costs and optimization difficulties, leading to significant performance degradation. As illustrated in Figure 6, the baseline model struggles with preserving global semantics and generating fine-grained visual details. In contrast, the proposed Semantics-Prompted DiT (SP-DiT) addresses these challenges, achieving significantly improved accuracy, for example, a 78% gain on the NYUv2 AbsRel metric. We further introduce a novel Cascaded DiT Design (Cas-DiT) that progressively increases the number of tokens. This coarse-to-fine design not only significantly improves efficiency, for example, reducing inference time by 30% on an RTX 4090 GPU, but also better models global context, leading to noticeable gains in accuracy.

### v2

@@ 2.1 Monocular Depth Estimation
Depth estimation can be broadly categorized into monocular yang2024depthv2 ; wang2025moge , stereo igev ; wang2024selective ; xu2023accurate ; igevpp ; guo2023openstereo ; cheng2025monster ; cheng2025monsterpp ; cheng2024adaptive , and sparse depth completion promptda methods. Early monocular depth estimation methods relied primarily on manually designed features saxena2008make3d ; hoiem2007recovering . The advent of neural networks revolutionized the field, though initial approaches eigen2014depth ; eigen2015predicting struggled with cross-dataset generalization. To address this limitation, scale-invariant and relative loss midas are introduced, enabling multi-dataset Li2018:CVPR ; blendedmvs ; diml ; hrwsi ; tartanair ; irs ; wsvd ; redweb ; hypersim ; liang2025sood++ training. Recent methods focus on improving the generalization ability yang2024depthv2 ; depthpro ; wang2025jasmine ; wang2025editor , depth consistency yang2024depthanyvideo ; chen2025video ; hu2024depthcrafter ; ke2024video , and metric scale bhat2023zoedepth ; li2024patchfusion ; li2024patchrefiner ; yin2023metric3d ; guizilini2023towards ; leres ; hu2024metric3dv2 ; piccinelli2024unidepth ; promptda of depth estimation. These methods converge towards using transformer-based architectures dpt . Concurrent works wang2025moge ; moge2 ; xu2025geometrycrafter explore point cloud representations to improve depth estimation performance. Several recent methods ji2023ddp ; duan2024diffusiondepth ; saxena2023monocular ; saxena2023surprising ; saxena2023zero ; zhao2023unleashing have attempted to use diffusion models for metric depth estimation. In contrast, our method focuses on relative depth and demonstrates improved generalization and fine-grained detail across a wide range of real-world scenes. Furthermore, our model significantly differs from these methods by introducing Semantics-Prompted DiT, which incorporates pretrained high-level semantic representations into the diffusion process, greatly enhancing performance.
@@ 2.2 Diffusion Generative Models
Diffusion generative models ddpm ; ddim ; dit ; yu2024representation ; yao2024fasterdit ; yao2025reconstruction ; zhu2025dig have demonstrated impressive results in image and video generation. Early approaches ddpm ; ho2022classifier ; ho2022cascaded such as DDPM ddpm operate directly in the pixel space, enabling high-fidelity generation but incurring significant computational costs, especially at high resolutions. To address this limitation, Latent Diffusion Models perform the diffusion process in a lower-dimensional latent space obtained via a VAE, as popularized by Stable Diffusion stablediff . This design significantly improves training and inference efficiency and has been widely adopted in recent works sd3 ; yao2025reconstruction ; yu2024representation ; zhu2024dig ; flux2024 ; podell2023sdxl ; yang2024cogvideox .
@@ 4.2 Ablations and Analysis
Component-wise ablation analysis. We adopt the vanilla DiT dit model as our baseline and conduct ablations on our proposed modules. Quantitative results are shown in Table 2. Directly performing diffusion generation in high-resolution pixel space is highly challenging due to substantial computational costs and optimization difficulties, leading to significant performance degradation. As illustrated in Figure 6, the baseline model struggles with preserving global semantics and generating fine-grained visual details. In contrast, the proposed Semantics-Prompted DiT (SP-DiT) addresses these challenges, achieving significantly improved accuracy, for example, a 78% gain on the NYUv2 AbsRel metric. We further introduce a novel Cascaded DiT Design (Cas-DiT) that progressively increases the number of tokens. This coarse-to-fine design not only significantly improves efficiency, for example, reducing inference time by 30% on an RTX 4090 GPU, but also better models global context, leading to noticeable gains in accuracy.